OpenStudy (patrickweewoos):

Which value for x makes the sentence true? 13x = 8x + 10

OpenStudy (patrickweewoos):

@phi

OpenStudy (igreen):

13x = 8x + 10 Subtract 8x to both sides: 5x = 10 Divide 10 / 5, can you do that? @PatrickWeeWoos

OpenStudy (patrickweewoos):

2

OpenStudy (igreen):

Yep, so x = 2.

OpenStudy (patrickweewoos):

Which value for x makes the sentence true? 13x - 4 = 6x + 24

OpenStudy (phi):

what do you get if you subtract 6x from both sides? in other words, take away 6 x's from each side what is left ?

OpenStudy (patrickweewoos):

24+13

OpenStudy (patrickweewoos):

its 23

OpenStudy (phi):

Let's see start with 13x - 4 = 6x + 24 write -6x on both sides to show we are taking 6 x's away 13x - 6x - 4 = 6x - 6x + 24 on the right side, do you know what we get ?

OpenStudy (patrickweewoos):

23

OpenStudy (phi):

what is 6x take away 6x ?

OpenStudy (patrickweewoos):

0

OpenStudy (phi):

so we can replace 6x - 6x with 0 in 3x - 6x - 4 = 6x - 6x + 24 to get 3x - 6x - 4 = 0 + 24 what is 0 + 24 ?

OpenStudy (phi):

I assume you know 0+24 simplifies to 24 so you now have 13x - 6x - 4 = 24 on the left side you have 13 x's and take away 6 x's how many x's are left ?

OpenStudy (patrickweewoos):

i got it wrong

OpenStudy (phi):

Does that mean you do not want to finish this problem ?

OpenStudy (phi):

If you have a bunch of these problems, you use the same idea to get the answer for each one. Once you get the idea, you can answer the questions.

OpenStudy (phi):

so you now have 13x - 6x - 4 = 24 on the left side you have 13 x's and take away 6 x's how many x's are left ?